OpenStudy (anonymous):

how do you solve for Y? x + 2y = -6.

OpenStudy (diyadiya):

Step1: subtract x from both sides

OpenStudy (diyadiya):

can you do that?

OpenStudy (anonymous):

i just barely learned these last semester but i need to refresh my mind on it i have it mixed up

OpenStudy (diyadiya):

x + 2y -x = -6 - x what will you get now?

OpenStudy (diyadiya):

we are trying to isolate y so we subtract x from both sides what do you get now?

OpenStudy (anonymous):

x+2y=-6 isn't it?

OpenStudy (diyadiya):

that's the question

OpenStudy (diyadiya):

now we subtract x on both sides

OpenStudy (anonymous):

i know but is that how its supposed to be simplified

OpenStudy (diyadiya):

Yes! we haven't completed thats just the first step ,we'll go stepwise

OpenStudy (anonymous):

ok so then which goes first

OpenStudy (diyadiya):

x + 2y = -6 we subtract x on both sides x + 2y -x = -6 - x now tell me what is x - x =?

OpenStudy (anonymous):

wouldnt it cancel

OpenStudy (diyadiya):

Yes! so what do you get?

OpenStudy (anonymous):

0

OpenStudy (diyadiya):

Yes